OpenStudy (anonymous):

find the product -8x(-3x^3 - 3x^2 - x -8)

OpenStudy (campbell_st):

do you know the index law from multiplication of the same base?

OpenStudy (anonymous):

no

OpenStudy (campbell_st):

ok... so lets start with a basic A letter on its own has a power of 1 \[x^1 = x\] then the law you need is when multiplying the same base, add the powers.. \[x^a \times x^b = x^{ a + b}\] does that make sense?

OpenStudy (campbell_st):

so you have \[(-8 \times -3) \times x^1 \times x^3 + (-8 \times -3) \times x^1 \times x^2 + (-8 \times -1) \times x^1 \times x^1 + (-8 \times -8) \times x^1\]

OpenStudy (anonymous):

Perfect! Thank you!